OpenStudy (kayders1997):

Using u sub find the integral of sin2x/1+cos^2x

OpenStudy (kayders1997):

@zepdrix

ganeshie8 (ganeshie8):

let 1+cos^2x = u

OpenStudy (kayders1997):

Okay

ganeshie8 (ganeshie8):

rest should be easy if you're familiar with the general method of substitution

OpenStudy (kayders1997):

Let me try it I'm not the best at it

OpenStudy (kayders1997):

I'll let you know what I get

ganeshie8 (ganeshie8):

sounds good

OpenStudy (kayders1997):

So wait when using u sub were trying to get sin2x right?

ganeshie8 (ganeshie8):

may i suggest you watch this awesome video on u substitution first https://www.khanacademy.org/math/integral-calculus/integration-techniques/u-substitution/v/u-substitution

OpenStudy (kayders1997):

Okay

OpenStudy (kayders1997):

Okay I watched it that was pretty helpful :)

OpenStudy (kayders1997):

Okay I got ln|1+cos^2|+c?

OpenStudy (kayders1997):

@ganeshie8 @zepdrix is it right?. :D

OpenStudy (kayders1997):

Wait...its -ln

zepdrix (zepdrix):

Ya that sounds better :o

OpenStudy (kayders1997):

The derivative of cosx is -sinx

OpenStudy (kayders1997):

Ughhhhhhh

OpenStudy (kayders1997):

How do you don't this one find the integral of x^2square root 2+x?

zepdrix (zepdrix):

This?\[\large\rm \int\limits x^2\sqrt{2+x}~dx\]There is a little trick to it.

OpenStudy (kayders1997):

Yeah I put 2+x as my u I don't think that was a wise move

zepdrix (zepdrix):

\[\large\rm u=2+x\]Now, before you go looking for your \(\large\rm du\), let's modify this to create an x^2.

zepdrix (zepdrix):

Subtracting 2,\[\large\rm u-2=x\]

OpenStudy (kayders1997):

Hmmm okay

zepdrix (zepdrix):

Squaring,\[\large\rm (u-2)^2=x^2\]

zepdrix (zepdrix):

To find your \(\large\rm du\), use the original substitution that we set up.

OpenStudy (kayders1997):

Oh wow

OpenStudy (kayders1997):

Wait what

OpenStudy (kayders1997):

Your saying du=dx use the original one?

zepdrix (zepdrix):

Good, that looks right. So, this was a little trickier, ya? We instead have `three pieces` to replace.

OpenStudy (kayders1997):

Yes

zepdrix (zepdrix):

\[\rm \color{orangered}{u=2+x}\]\[\rm \color{royalblue}{(u-2)^2=x^2}\]\[\rm \color{green}{du=dx}\]Into our integral,\[\large\rm \int\limits \color{royalblue}{x^2}\sqrt{\color{orangered}{2+x}}~\color{green}{dx}\]

OpenStudy (kayders1997):

Woooo let me think this lol

OpenStudy (kayders1997):

So (u-2)^2times square root u du?

zepdrix (zepdrix):

\[\large\rm \int\limits (u-2)^2u^{1/2}du\]Mmm k good.

zepdrix (zepdrix):

No nice tricks from this point, just expand out the square, multiply the u^{1/2} into each, and integrate term by term, power rule.

OpenStudy (kayders1997):

Okay

OpenStudy (kayders1997):

Okay

OpenStudy (kayders1997):

U^5/2-2u^3/2+4u^1/2

zepdrix (zepdrix):

woops, middle term is -4u^{3/2}

OpenStudy (kayders1997):

oh yeah oops :D

OpenStudy (kayders1997):

I was nervous I was like wait am I supposed to multiply exponents I always forget... not a good thing, so basic things I should probably be reviewing

OpenStudy (kayders1997):

So now take the integral of it you get

OpenStudy (kayders1997):

it wont let me use the equation thingy...dumb

zepdrix (zepdrix):

twiddla? :d lol

OpenStudy (kayders1997):

what about it? you wanna go on there?

zepdrix (zepdrix):

ya maybe :D if it wont let you use the equation tool

OpenStudy (kayders1997):

yeah lets do it

zepdrix (zepdrix):

https://www.twiddla.com/mvbswa